# LeGrad: An Explainability Method for Vision Transformers via Feature Formation Sensitivity

### v1

@@ Abstract
Vision Transformers (ViTs), with their ability to model long-range dependencies through self-attention mechanisms, have become a standard architecture in computer vision. However, the interpretability of these models remains a challenge. To address this, we propose LeGrad, an explainability method specifically designed for ViTs. LeGrad computes the gradient with respect to the attention maps of ViT layers, considering the gradient itself as the explainability signal.We aggregate the signal over all layers, combining the activations of the last as well as intermediate tokens to produce the merged explainability map. This makes LeGrad a conceptually simple and an easy-to-implement tool for enhancing the transparency of ViTs. We evaluate LeGrad in challenging segmentation, perturbation, and open-vocabulary settings, showcasing its versatility compared to other SotA explainability methods demonstrating its superior spatial fidelity and robustness to perturbations. [^1] A demo and the code is available at https://github.com/WalBouss/LeGrad
Figure 1: LeGrad explainability maps: For a given vision-language model and a textual prompt, LeGrad generates a heatmap indicating the part of the image that is most sensitive to that prompt. LeGrad performs equally well across a large range of model sizes, ranging from ViT-B/16(150M params.) to ViT-bigG/14(2B params.).
@@ 1 Introduction
In the realm of computer vision, Vision Transformers (ViTs)[16] have revolutionized the field with their ability to model long-range dependencies through self-attention mechanisms [43, 27, 45, 4, 30]. Explanability methods designed for convolutional or feed-forward neural networks are not directly applicable to ViTs due to their architectural requirements, like GradCAM’s [37] reliance on convolutional layers and Layer-wise Relevance Propagation’s (LRP) [5] specific layer-wise propagation rules. While ViT-specific interpretability techniques exist, including adaptations of traditional methods [12, 37, 13], attention-based techniques [1, 44, 12, 13], and text-based explanations [22, 20, 1], the interpretability especially of large-scale architectures remains a challenge.
To address this, we propose a Layerwise Explainability method that considers the Gradient with respect to the attention maps, LeGrad. LeGrad is specifically designed for ViTs as leverages the self-attention mechanism intrinsic to ViTs to generate relevancy maps highlighting the most influential parts of an image for the model’s prediction. Compared to other methods, LeGrad computes gradient with respect to the attention maps, as e.g. opposed to CheferCam [12, 13], which uses the gradient to weight the attention maps. This is done independently for each layer. The final explainability signal is then pooled over all layers of the ViT. Note that using a layerweise gradient, compared to other signals, allows to sum up over different layers without further need for normalization. To further improve the signal, the gradient is clipped by a ReLU function preventing negative gradients to impact positive activations. The proposed approach is conceptually simple and versatile, as it only requires the gradient w.r.t. to the ViT’s attention maps. This facilitates its adoption across various applications and architecture, including larger ViTs such as ViT-BigG as well as attention pooling architectures [26, 49], enhancing the transparency of ViTs.
We evaluate the proposed method for various ViT backbones on three challenging tasks: segmentation, open-vocabulary detection, and perturbation on ImageNet [34, 19] as well as OpenImagesV7 [6]. It shows that while current methods struggle especially with the diverse object categories in OpenImagesV7, LeGrad can achieve performance gains of 2×2$\times$2 ×5×5$\times$5 ×, reaching a score of $48.4$ p-mIoU on OpenImagesV7 using OpenCLIP-ViT-B/16. Furthermore, we demonstrate the applicability of LeGrad to very large models, such as the ViT-BigG/14 [14] with 2.5 billion parameters. Finally, it also adapts well to different feature aggregation strategies employed by ViTs [32, 49], making it a versatile tool for explainability.
@@ 2 Related Work
Gradient-Based Explanation Methods Feature-attribution methods (also known as saliency methods) are a commonly used explanation technique that explain model decisions by assigning a score to each image pixel representing its importance to the model’s output. Largely, these methods can be categorized into two groups [29] — gradient-based methods that compute explanations based on the gradient of the model’s prediction with respect to each input pixel [38, 17, 41, 40, 39, 24, 37] and perturbation-based methods that measure pixel-importance by successively perturbing the input images and measuring the impact on the model’s output [28, 33, 31, 11, 48]. While both types of methods have been used successfully to identify reasoning, spurious correlation, and trustworthiness in traditional computer vision models [7, 10], gradient-based methods are often more computationally efficient since they only require a single backwards pass and simple to understand since they are a direct function of the model’s parameters and do not rely on additional models or image modifications. However, many existing gradient-based methods were designed for convolutional and feed-forward model architectures, so it is non-trivial to directly apply them to ViTs since ViTs do not contain spatial feature maps and include complex interactions between patches induced by the self-attention mechanism. Instead, we build on the simplicity and computational efficiency of gradient-based explanation methods to develop LeGrad, a gradient-based feature-attribution method specifically designed for ViT architectures.
While most gradient-based methods were designed prior to the widespread use of ViTs, researchers have recently made efforts to adapt existing methods or develop new ones specifically for transformers. Chefer et al. [12] extend LRP [5] to transformers by integrating gradients within the self-attention layers. However, this approach is computationally heavy and is inflexible to architecture changes as it requires specific implementation for each module of the network. To circumvent that complexity, CheferCAM [13] weights the attention by their gradient and aggregates it through the layer via matrix multiplication. However, the use of gradients to weigh the attention heads’ importance makes this method class-specific. LeGrad on the other hand is class-agnostic and computationally efficient, making it simple to use and more effective than these other methods (see Sec. 4 for a quantitative evaluation).
Attention-Based Explanation Methods A separate line of research has proposed using attention maps, as opposed to gradients, as a way to explain for transformers’ decisions [1, 44]. One attention-based method, rollout [1], traces the flow of importance through the transformer’s layers by linearly combining the attention maps via matrix multiplication. Attention flow [1] contextualizes the attention mechanism as a max-flow problem; however, it is computationally demanding and has not been extensively evaluated for vision tasks. While these methods offer insights into the attention mechanism, they often neglect the non-linear interactions between attention heads and the subsequent layers. Moreover, they may not adequately distinguish between positive and negative contributions to the final decision, leading to potentially misleading interpretations, as found in Chefer et al. [12]. While LeGrad also focuses on the attention maps to generate its explanation, it uses the gradient w.r.t. to the attention maps, thereby assessing the sensitivity of the attention maps to a change in the patch tokens. LeGrad does not attempt to track the flow of information throughout the transformer layers but, rather, focuses on assessing the sensitivity of the model prediction at each layer for every location of the image.
Text-based Explanations of Intermediate Representations Research has also explored text-based methods for interpreting intermediate representations in vision models. For instance, leveraging CLIP’s language-image space, researchers have provided text descriptions for active neuron regions [22, 20] and projected model features into text-based concept banks [18]. In particular, TextSpan [18] focuses on the explainability of CLIP-like models. It refrains from using gradient computation by aggregating the intermediate features’ similarities along a given text direction, creating a relevancy map for a text query. LeGrad advances this line of work by focusing on the sensitivity of feature representations within ViTs, offering a clear and concise method to generate relevancy maps that can be adapted to various feature aggregation strategies.
Evaluating Explanation Methods As the popularity of feature-attribution based explanation methods has grown, so has research evaluating their ability to faithfully reflect the model’s decision making process [41, 46, 25, 35, 23, 3, 2, 42, 21, 50]. Faithfulness is an important criteria because explanation methods that highlight human-salient pixels may look correct even if the model relies on spurious correlations and give human interpreters unwarranted confidence in the model. These evaluations can be categorized based on the aspect of faithfulness that they test [8], such as the method’s sensitivity to meaningful changes to the input, model, or data [41, 46, 25, 35, 23, 3, 2, 42] and its ability to highlight a succinct and meaningful pixel region [21, 50]. We evaluate LeGrad on both of these types of tests, demonstrating that it is faithful to the model’s decision making process by testing that removing LeGrad’s important pixels impacts the model’s output and checking that LeGrad’s output corresponds to a concise object region.
@@
Figure 2: Overview of LeGrad: Given a text prompt or a classifier $\mathcal{C}$, for each layer $l$, an activation $s^{l}$ is computed and used to compute the explainability of that layer. The layerwise explainability maps are then merged to produce LeGrad’s output.
@@ 3 Method
3.1 Feature formation in ViTs
@@
$\hat{Z}^{l}$ $=\text{MSA}^{l}(Z^{l-1})+Z^{l-1},$ (1)
$Z^{l}$ $=\text{MLP}^{l}(\hat{Z}^{l})+\hat{Z}^{l}.$
After $L$ layers, the image representation can be obtained via various strategies:
[CLS] token: The class token approach, as introduced in the original ViT paper [16], uses the processed class token as the image embedding $\bar{z}_{[CLS]}=z^{L}_{0}$. This method relies on the transformer’s ability to aggregate information from the patch tokens into the class token during training.
Attentional Pooler: as used in [26, 47, 49], employs an additional multi-head attention layer with a learnable query token $q_{pool}\in\mathbb{R}^{d}$. This token interacts with the final layer patch tokens to produce the pooled representation $\bar{z}_{AttnPool}$:
@@
Independent of the feature aggregation strategy, it is important for an explainability method to account for the iterative nature of feature formation in ViTs. The goal should be to capture the contributions of all layers for the final representation, providing insights into the model’s decision-making process.
@@ 3.2 Explainability Method: LeGrad
We denote the output tokens of each block $l$ of the ViT as $Z^{l}=\{z_{0}^{l},z_{1}^{l},\dots,z_{n}^{l}\}\in\mathbb{R}^{(n+1)\times d}$, where $d$ is the dimensionality of each token.
Consider a classifier $\mathcal{C}\in\mathbb{R}^{d\times C}$, where $C$ is the number of classes. This classifier can be learned during training for a supervised classification task (e.g., ImageNet) or can be formed from text embeddings of class descriptions for a zero-shot classifier, as in the case of vision-language models like CLIP, generating a prediction $\bar{y}$. This prediction $\bar{y}$ is obtained by passing the aggregated feature representation of the ViT, noted $\bar{z}$, through the classifier $\mathcal{C}$ and selecting the class with the highest activation:
$\bar{y}=\bar{z}\cdot\mathcal{C}\in\mathbb{R}^{C},$ (3)
The following is the description of how to obtain the explainability map for a single layer, using the last layer as an example and then how to generalize it to multiple layers. The overall method is visualized Figure 2.
@@
Process for a Single Layer: To compute a 2D map that highlights the image regions most influential for the model’s prediction of a particular class, we focus on the activation with respect to the target class $\hat{c}$, denoted by $s=\bar{y}_{[\hat{c}]}$. The attention operation within a ViT is key to information sharing, and thus our method concentrates on this process.
We compute the gradient of the activation $s$ with respect to the attention map of the last layer, as shown in Figure 3, denoted as $\mathbf{A}^{L}\in\mathbb{R}^{h\times n\times n}$:
@@
where $h$ is the number of heads in the self-attention operation. Negative contributions are eliminated by applying a ReLU function, and the gradient is averaged across the patch and head dimensions:
@@
The final explainability map is obtained by removing the column corresponding to the [CLS] token to isolate the influence of the patch tokens, reshaping it into a 2D map, and applying min-max normalization:
@@
Prior to normalization, we average the explainability maps from each layer:
@@
$\mathbf{E}$ $=\text{norm}(\text{reshape}(\bar{\mathbf{E}}))\in R^{W\times H}.$
This approach queries each patch token at a given layer about its influence on the prediction at that stage.
@@ 4 Experiments
4.1 Evaluation Settings
Implementation details: For the empirical evaluation of LeGrad, we utilized a suite of Vision Transformer models sourced from the OpenCLIP library [14], trained on the LAION dataset [36]. The models employed in our experiments include ViT-B/16, ViT-L/14, ViT-H/14, and ViT-bigG/14.For SigLIP [49], we use the ViT-B/16 with the "webli" weights. Each model processes input images at a resolution of $224\times 224$ pixels. A key advantage of LeGrad is its hyperparameter-free nature, which obviates the need for hyperparameter tuning. Furthermore, we do not retraining or fine-tune the vision-language models, thereby demonstrating their capabilities to localize queries by solely being trained on image-level data without any localization annotations during training or fine-tuning phases.
@@ 4.1 Evaluation Settings
[…]
4.2 Comparison to State-of-the-Art
Segmentation Analysis: We first evaluate LeGrad as well as other methods in the context of image segmentation using the ImageNet-segmentation dataset. As shown in Table 2, LeGrad achieved a mIoU of 58.7%, surpassing other SOTA explainability methods. Notably, it outperformed CheferCAM as a gradient-based method for ViTs, and TextSpan as a non-gradient-based method, indicating its robustness in capturing relevant image features for classification tasks.
Open-Vocabulary Performance: We further evaluate the performance on the OpenImagesV7 dataset testing the capabilities of all methods in handling diverse object categories. Note that for GradCAM we searched over the layers and took the one that was performing the best. As shown in Table 2, LeGrad outperforms all other SOTA methods, with performance gains ranging from 2×2$\times$2 × to 5×5$\times$5 × compared to the second-best performing method. This can be seen as an indicator for LeGrad’s capacity for fine-grained recognition.
Perturbation-Based Evaluation: We further performed the perturbation-based evaluation on the ImageNet-val datasets. As detailed in Table 3, LeGrad’s performance is here comparable to TextSpan for positive perturbations and slightly superior for negative perturbations. For all other methods, LeGrad outperformes both attention-based (e.g., "rollout" and "raw attention") and gradient-based methods (e.g., GradCAM, AttentionCAM, and CheferCAM) across various model sizes and for both predicted and ground truth classes, emphasizing its ability to identify and preserve critical image regions for accurate classification.
Performance on SigLIP: We finally evaluate the adaptability b regarding the performance on SigLIP-B/16, a Vision-Language model employing an attentional pooler as shown in Table 4. The results underscore the methods performance across both negative and positive perturbation-based benchmarks. Notably, in the open-vocabulary benchmark on OpenImagesV7, LeGrad achieved a p-mIoU of 25.4, significantly surpassing GradCAM’s 7.0 p-mIoU, the next best method. These findings affirm the versatility of LeGrad, demonstrating its robust applicability to various pooling mechanisms within Vision Transformers. Further details on the methodological adaptations of LeGrad and other evaluated methods for compatibility with SigLIP are provided in the annex.
@@ 4.2 Comparison to State-of-the-Art
Table 1: Object Segmentation: method comparison on ImageNet-S using a CLIP(ViT-B/16) model.
@@
Method B/16 L/14 H/14 bigG/14
rollout 8.75 6.85 5.82 4.74
Raw attention 0.94 1.60 0.85 0.07
GradCAM 8.72 2.80 2.46 0.14
AttentionCAM 5.87 4.74 1.20 0.72
CheferCAM 5.87 2.51 9.49 9.71
TextSpan 9.44 21.73 23.74 20.18
LeGrad 48.38 47.69 46.51 36.47
Table 2: Open-Vocabulary Segmentation Performance: methods comparison on OpenImagesV7 using different openCLIP model size.

### v2

@@ Abstract
Vision Transformers (ViTs) have become a standard architecture in computer vision. However, because of their modeling of long-range dependencies through self-attention mechanisms, the explainability of these models remains a challenge. To address this, we propose LeGrad, an explainability method specifically designed for ViTs. LeGrad computes the gradient with respect to the attention maps of single ViT layers, considering the gradient itself as the explainability signal. We aggregate the signal over all layers, combining the activations of the last as well as intermediate tokens to produce the merged explainability map. This makes LeGrad a conceptually simple and an easy-to-implement method to enhance the transparency of ViTs. We evaluate LeGrad in various setups, including segmentation, perturbation, and open-vocabulary settings, showcasing its improved spatial fidelity as well as its versatility compared to other SotA explainability methods.
Figure 1: LeGrad explainability maps: For a given VLM and an input textual prompt, LeGrad generates a heatmap indicating the part of the image that is most sensitive to that prompt. Examples shown for OpenCLIP ViT-B/16(150M params.) and ViT-bigG/14(2B params.).
@@ 1 Introduction
Vision Transformers (ViTs)[12] have significantly influcened the field of computer vision with their ability to model long-range dependencies through self-attention mechanisms. But explanability methods designed for convolutional or feed-forward neural networks are not directly applicable to ViTs due to their architectural requirements, like GradCAM’s [32] reliance on convolutional layers and Layer-wise Relevance Propagation’s (LRP) [2] specific layer-wise propagation rules. While ViT-specific explainability techniques exist, including adaptations of traditional methods [8, 32, 9], attention-based techniques [1, 37, 8, 9], and text-based explanations [21, 17, 1], the explainability those architectures remains a challenge.
To address this problem, we propose LeGrad, a Layerwise Explainability method that considers the Gradient with respect to the attention maps. LeGrad is specifically designed for ViTs as it leverages the self-attention mechanism to generate relevancy maps highlighting the most influential parts of an image for the model’s prediction. Compared to other methods, LeGrad uses the gradient with respect to the attention maps as the explanatory signal, as e.g. opposed to CheferCam [8, 9], which uses the gradient to weight the attention maps. This is done independently for each layer. The final explainability signal is then pooled over all layers of the ViT. Note that using a layerwise gradient, compared to other signals, allows to sum up over different layers without further need for normalization. To further improve the signal, the gradient is clipped by a ReLU function preventing negative gradients to impact positive activations (see Figure 2 for details). The approach is conceptually simple and versatile, as it only requires the gradient w.r.t. to the ViT’s attention maps. This facilitates its adoption across various applications and architecture, including larger ViTs such as ViT-BigG as well as attention pooling architectures [24, 41].
We evaluate the proposed method for various ViT backbones on four challenging tasks, segmentation, open-vocabulary detection, perturbation, and audio localization, spanning over various datasets, incl. ImageNet [31, 15], OpenImagesV7 [3], and ADE20KSound/SpeechPrompted [19]. It shows that while current methods struggle with the diverse object categories in OpenImagesV7, LeGrad reaches a score of $48.4$ p-mIoU on OpenImagesV7 using OpenCLIP-ViT-B/16. Furthermore, we demonstrate the applicability of LeGrad to very large models, such as the ViT-BigG/14 [10] with 2.5 billion parameters while also adapting well to different feature aggregation strategies employed e.g. by SigLIP [41]. Finally, LeGrad also establishes a new SoTA on zero-shot sound localization on ADE20KSoundPrompted scoring $+14mIoU$ over previous SoTA.
@@ 2 Related Work
Gradient-Based Explanation Methods Feature-attribution methods are a commonly used explanation technique that explains model decisions by assigning a score to each image pixel, representing its importance to the model’s output. Generally, these methods can be categorized into two groups [26] — gradient-based methods that compute explanations based on the gradient of the prediction of the model with respect to each input pixel [33, 13, 36, 35, 34, 23, 32] and perturbation-based methods that measure pixel importance by successively perturbing the input images and measuring the impact on the model output [25, 30, 28, 6, 40]. While both types of methods have been used successfully to identify correlations and trustworthiness in traditional computer vision models [4, 7], gradient-based methods are often more computationally efficient since they only require a single backwards pass. Further, they are easy to interpret since they are a direct function of the model’s parameters and do not rely on additional models or image modifications. However, many existing gradient-based methods were designed for convolutional and feed-forward architectures, so it is non-trivial to directly apply them to ViTs since ViTs do not contain spatial feature maps and include complex interactions between patches induced by the self-attention mechanism.
@@
Figure 2: Overview of LeGrad: Given a text prompt or a classifier $\mathcal{C}$, an activation $s^{l}$ is computed for each layer $l$. The activation $s^{l}$ is then used to compute the explainability map of that layer. The layerwise explainability maps are then merged to produce LeGrad’s output.
As most gradient-based methods were designed prior to the widespread use of ViTs, researchers have recently made efforts to adapt existing methods and to develop new ones specifically for transformers. Chefer et al. [8] extend LRP [2] to transformers by integrating gradients within the self-attention layers. However, this approach is computationally heavy and is inflexible to architecture changes as it requires specific implementation for each module of the network. To circumvent that complexity, CheferCAM [9] weights the attention by their gradient and aggregates it through the layer via matrix multiplication. However, the use of gradients to weigh the attention heads’ importance makes this method class-specific.
Explanability Methods for ViT A separate line of research has proposed using ViT attention maps, as opposed to gradients, as a way to explain for transformers’ decisions [1, 37]. One attention-based method, rollout [1], traces the flow of importance through the transformer’s layers by linearly combining the attention maps via matrix multiplication. Attention flow [1] contextualizes the attention mechanism as a max-flow problem; however, it is computationally demanding and has not been extensively evaluated for vision tasks. While these methods offer insights into the attention mechanism, they often neglect the non-linear interactions between attention heads and the subsequent layers. Moreover, they may not adequately distinguish between positive and negative contributions to the final decision, leading to potentially misleading interpretations, as found in Chefer et al. [8]. Compared to that LeGrad uses the gradient w.r.t. to the attention maps, thereby assessing the sensitivity of the attention maps to a change in the patch tokens.
Vision-Language Explainability Methods Research has also explored vision-langauge models for interpreting representations in vision models. For instance, leveraging CLIP’s language-image space, researchers have provided text descriptions for active neuron regions [21, 17] and projected model features into text-based concept banks [14]. In particular, TextSpan [14] focuses on the explainability of CLIP-like models. It refrains from using gradient computation by aggregating the intermediate features’ similarities along a given text direction, creating a relevancy map for a text query. LeGrad advances this line of work by focusing on the sensitivity of feature representations within ViTs to generate relevancy maps that can be adapted to various feature aggregation strategies.
@@ 3 Method
3.1 Background: Feature formation in ViTs
@@
$\hat{Z}^{l}=\text{MSA}^{l}(Z^{l-1})+Z^{l-1},\quad Z^{l}=\text{MLP}^{l}(\hat{Z}^{l})+\hat{Z}^{l}.$ (1)
After $L$ layers, the visual features can be obtained via:
[CLS] token: The class token approach, as introduced in ViT [12], uses the processed class token as the image embedding $\bar{z}_{[CLS]}=z^{L}_{0}$. This method relies on the transformer’s ability to aggregate information from the patch tokens into the class token during training.
Attentional Pooler: Attention Pooling, as e.g. used in SigLIP [41] employs an multi-head attention layer [24, 39] with a learnable query token $q_{pool}\in\mathbb{R}^{d}$. This token interacts with the final layer patch tokens to produce the pooled representation $\bar{z}_{AttnPool}$:
@@
Independent of the feature aggregation strategy, it is important for an explainability method to account for the iterative nature of feature formation in ViTs and to capture the contributions of all layers for the final representation. LeGrad addresses this by fusing information from each layer, allowing for both, a granular as well as a joint, holistic interpretation of the model’s predictions.
@@ 3.2 Explainability Method: LeGrad
We denote the output tokens of each block $l$ of the ViT as $Z^{l}=\{z_{0}^{l},z_{1}^{l},\dots,z_{n}^{l}\}\in\mathbb{R}^{(n+1)\times d}$, where $d$ is the dimensionality of each token, and $\bar{z}^{l}$ is the average over the tokens of the respective layer $l$, defined as $\bar{z}^{l}=\frac{1}{n+1}\sum_{i=0}^{n}z_{i}^{l}$
Consider a mapping model $\mathcal{C}\in\mathbb{R}^{d\times C}$, that maps the token dimension $d$ to $C$ logits. This classifier can be learned during training for a supervised classification task (e.g., ImageNet) and in that case $C$ is the number of classes. Or it can be formed from text embeddings of some prompts in the case of zero-shot classifier, e.g., vision-language models like CLIP, then $C$ is the number of prompts. For a given layer $l$, the mapping model $\mathcal{C}$ then generates a prediction $\bar{y}^{l}$, which can be the output of the classifier or, in case of vision-language models, the vector of results of dot products of the text and image embeddings. This prediction $\bar{y}^{l}$ is obtained by passing the aggregated feature representation of the ViT, noted $\bar{z}^{l}$, through the mapping $\mathcal{C}$:
$\bar{y}^{l}=\bar{z}^{l}\cdot\mathcal{C}\in\mathbb{R}^{C}.\vspace{-0.5em}$ (3)
Note that most explainability methods only use the final outputs of the model. We argue that also leveraging intermediate representations is beneficial (see Section 4.7). The following we first describe how to obtain the explainability map for a single layer, using an arbitrary layer $l$ as an example and then generalize it to multiple layers. The overall method is visualized Figure 2.
@@
Process for a Single Layer: To compute a 2D map that highlights the image regions most influential for the model’s prediction of a particular class, we focus on the activation with respect to the target class/prompt $\hat{c}$, denoted by $s^{l}=\bar{y}^{l}_{[\hat{c}]}$. The attention operation within a ViT is key to information sharing, and thus our method concentrates on this process. We compute the gradient of the activation $s^{l}$ with respect to the attention map of layer $l$, as shown in Figure 3, denoted as $\mathbf{A}^{l}\in\mathbb{R}^{h\times n\times n}$:
@@
where $h$ is the number of heads in the self-attention operation. Negative gradients are discarded by a ReLU function (noted $(.)^{+}$), and the gradient is averaged across the patch and head dimensions:
@@
To obtain the final explainability map, the column corresponding to the [CLS] token is removed, only considering the patch tokens, reshaped into a 2D map, and a min-max normalization is applied:
@@
We then average the explainability maps from each layer:
@@
And finally we reshape to the original image size and apply min-max normalization:
$\mathbf{E}$ $=\text{norm}(\text{reshape}(\bar{\mathbf{E}}))\in R^{W\times H}.$ (9)
This queries each patch token at a given layer about its influence on the prediction at that stage.
@@ 4 Experiments
4.1 Object Segmentation
@@
Table 3 compares LeGrad as well as other methods in the context of image segmentation using the ImageNet-segmentation dataset. LeGrad achieved a mIoU of 58.7%, surpassing other SOTA explainability methods. Notably, it outperformed CheferCAM as a gradient-based method for ViTs, and TextSpan as a non-gradient-based method, indicating its robustness in capturing relevant image features for classification tasks.
@@
Table 1: Object Segmentation: method on ImageNet-S using an OpenCLIP(ViT-B/16) model trained on Laion2B.
@@
Method B/16 L/14 H/14
rollout 8.75 6.85 5.82
Raw attention 0.94 1.60 0.85
GradCAM 8.72 2.80 2.46
AttentionCAM 5.87 4.74 1.20
CheferCAM 5.87 2.51 9.49
TextSpan 9.44 21.73 23.74
LeGrad 48.38 47.69 46.51
Table 2: Open-Vocabulary Segmentation: methods comparison on OpenImagesV7 using different OpenCLIP model sizes.
Method fps
AttentionCAM $103$
GradCAM $108$
CheferCAM $21$
LRP $4.0$
TextSpan $3.8$
LeGrad $96$
Table 3: Inference speed comparison for ViT-B/16.
4.2 Open-Vocabulary localization
[…]
Table 3 evaluates the performance on OpenImagesV7, testing the capabilities of all methods in handling diverse object categories. Note that for GradCAM we searched over the layers and took the one that was performing the best. LeGrad outperforms all other SOTA methods, with performance gains ranging from 2×2$\times$2 × to 5×5$\times$5 × compared to the second-best performing method. This can be seen as an indicator for LeGrad’s capacity for fine-grained recognition.
ImageNet OpenImagesV7